OpenStudy (anonymous):

At a temparature of 20 degree C the common amoeba reproduces by splitting in half every 24 hours.if we start with a single amoeba how many will there be after a)8 days b)16 days?

OpenStudy (anonymous):

where are you finding the problem to solve it..?? i'll try to correct you..:)

OpenStudy (anonymous):

well i am giving SAT and from D.Rayners book :)@RaGhavv

OpenStudy (anonymous):

no no...tell me where are you stuck up??...i mean how you solved it...or not able to..i will correct you..!

OpenStudy (anonymous):

WELL i don't know :/

OpenStudy (jack1):

y(t) = a × e^kt Where: y(t) = value at time "t" a = value at the start k = rate of growth (when >0) or decay (when <0) t = time

OpenStudy (jack1):

this is an exponential growth rate question, can you work it out from the above formula...?

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

see..its quite easy..you actually can solve it by your logic,,,,not many equations are required to solve this...so you must first try...

OpenStudy (jack1):

y(t) = a × e^kt a = 1 amoeba population doubles every day, so after 1 day... population =2 y(t) = a × e^kt 2 = 1 x e^kt ... 1 day so t = 1 2 = 1 x e^1k 2 = e^k k = log 2 so now you have all the constants so when t = 8days y(t) = a × e^kt y = 1 × e^(log2)t y = 1 × e^(log2)×8 y = e^5.55(...ish) y = 256 so population after 8 days = 256

OpenStudy (zzr0ck3r):

i.e 2^8

OpenStudy (jack1):

@zzr0ck3r totally, but that only works if the rate is a simple number... ie doubles every 3 hrs...

OpenStudy (zzr0ck3r):

you said yourself it doubles every day:) \[2^{number-of-days}\]

OpenStudy (zzr0ck3r):

correct